# Artificial Inductive Bias for Synthetic Tabular Data Generation in Data-Scarce Scenarios

### v1

@@ 2.2.2 DRS
Figure 5: Block diagram for the DRS case. The inductive bias is introduced by training a DGM $p_{\theta}$ on $x_{r}$ using $S$ different seeds. The synthetic dataset $x_{g}$ contains data generated by each seed and serves as input to DRS. The starting point to fine-tune using the real data $x_{r}$ and obtain $p_{\hat{\theta}}$ is the DRS solution obtained, $\theta_{DRS}$.
Although MAML offers the potential to leverage the underlying structure of learning problems through a powerful optimization framework, it introduces a significant computational cost. Therefore, while we should seek for a trade-off between accuracy and computational efficiency, there is no approach for managing this trade-off. It needs an understanding of the domain-specific characteristics inherent to the meta-problem itself.
DRS presents an alternative meta-learning approach that circumvents the computational burden associated with bilevel optimization problems. Unlike MAML, DRS trains a model on the combined data from all tasks. This eliminates the need for the complex optimization procedures present in MAML, leading to a more computationally efficient solution. However, it is important to acknowledge that DRS offers an approximation to the ideal solution [49].
@@
In this context, $\mathcal{D^{*}}$ represents the aggregated synthetic data collection, $x_{g}$, obtained across all training seeds $S$. We refer to this approach as Domain-Randomized Search due to its alignment with the domain randomization method presented in [50] and its core principle of directly searching over a distribution of domains (tasks).
Figure 5 shows the application of the DRS procedure within the framework of our proposed methodology. Similar to the MAML case, $\theta_{DRS}$ serves as the initialization weights that we aim to identify $\theta_{0}$.
Both MAML and DRS offer complementary approaches with a trade-off between modeling complexity and optimization cost [49]. DRS delivers an approximate solution with lower computational demands, while MAML offers higher precision at the expense of greater computational resources. DRS is also advantageous when dealing with a limited number of learning tasks. In our case, where data generated by each seed ($s=1,2,...,S$) is considered a task, and $S$ typically takes values around $10$, DRS is expected to provide better solutions than MAML, aligning with the findings of [49]. Finally, note that DRS is similar to the pre-training approach. While both techniques aim to improve model performance, they utilize data differently. Pre-training leverages data from the single best VAE seed, whereas DRS capitalizes on data from all VAE seeds. This distinction reflects the core principle of DRS: exploring a wider range of possibilities by searching across a distribution of domains (tasks) represented by the various seed variations.
@@ 3.1 Data
The experiments were carried out on four public datasets obtained from the SDV environment [51], which also implements various data generation models, including the CTGAN implementation we use. The experiment design prioritized datasets with a sufficient number of samples. This allows us to create multiple data splits for various configurations of the different training and validation parameters. This approach comprehensively evaluates the proposed method under different parameter settings.
  • Adult: The Adult Census Income dataset [52] is a mixed-data dataset extracted from the 1994 U.S. Census. It comprises $32,561$ data points, each described by 14 features that encompass integer, categorical, and binary values. The dataset is used to predict whether an individual’s annual income exceeds \$50,000. It should be noted that the data set incorporates 13% missing values, concentrated exclusively within two specific variables: “workclass" and “occupation".
  • News: The News Popularity Prediction dataset [53] consists of 39,644 samples and 58 columns containing information about articles published on the Mashable news blog over two years. The objective is to predict the popularity of an article, measured by the number of social media shares. The dataset is multivariate, including both continuous and categorical variables.
  • King [^1] https://www.kaggle.com/datasets/harlfoxem/housesalesprediction: The King County House Sales dataset is a regression dataset containing 21,613 house sale prices for King County, Washington, including Seattle. It includes homes sold between May 2014 and May 2015. The 20 features provide multivariate information in numerical (integer and decimal) and categorical forms.
  • Intrusion: The KDD Cup 1999 Data [54] was used for The Third Knowledge Discovery and Data Mining Competition. It comprises 494,021 samples and 39 information characteristics used to classify connections in a military network environment. The dataset contains categorical, integer, and binary features.
This selection of datasets with varying sample sizes, feature dimensions, and domains allows us to assess the performance and generalizability of the proposed method in diverse data scenarios. As will be demonstrated in the following sections, the proposed method is effective in generating synthetic data that retain the key characteristics of the original datasets across a wide range of applications.
@@ 3.2 Validation metrics
To rigorously evaluate the effectiveness of our proposed method in capturing the real data distribution, we strictly adhere to the validation approach described in [20]. This approach leverages a probabilistic classifier (discriminator) to estimate the ratio of probability densities between the real and synthetic distributions, subsequently calculating the KL and JS divergences. Traditional validation methods often focus on individual data points and the marginal distribution of each separate feature. In contrast, this approach considers the entire data distribution, including complex relationships between features. Additionally, divergences are robust to noise and offer clear interpretations, making them ideal for evaluating the DGM’s effectiveness. This provides a comprehensive approach to measuring the discrepancy between two probability distributions, making them suitable for assessing the similarity between real data $p(x_{r})$ and the distribution of the synthetic data generated by the DGM $p_{\theta}$.
@@
The discriminator network plays a crucial role in the validation process. This neural network architecture is trained to distinguish between real and synthetic data samples. The network receives two sets of samples as input. The first consists of $M$ samples from the real data distribution $p(x_{r})$ labeled as class $1$, and the second consists of $M$ samples labeled as class 00 from the synthetic data distribution generated by the DGM $p_{\theta}$ or $p_{\hat{\theta}}$, depending on the $N$ number of samples in the dataset. During training, the discriminator aims to learn a decision boundary that effectively separates these two sets of samples. This process forces the discriminator to capture the underlying differences between the real and synthetic distributions. Once the discriminator network is trained, it is used to estimate the KL and JS divergences between the real and synthetic probability distributions. This estimation involves using $L$ samples from each distribution and feeding them to the trained discriminator. The output probabilities of the discriminator for these samples are then used to compute the KL and JS divergence metrics.
Figure 6 illustrates the overall scheme of the approach. The figure emphasizes the separate but related components: the inductive bias generator for synthetic data creation and the validation process with the discriminator network. It also highlights the number of samples used from each distribution for each step of the process ($N$ to generate samples, $M$ to train the discriminator, and $L$ to estimate divergence). By adhering to this rigorous validation approach, we ensure that our proposed methodology for generating synthetic data from small datasets is thoroughly evaluated and its effectiveness is quantitatively demonstrated using established metrics like KL and JS divergences. Note that, as stated in [20], $M$ and $L$ need to be large enough to prevent inaccurate divergence estimations, so we consider this in our experiments.
@@
Figure 6: General Scheme of the proposed approach. Overall scheme of the approach and its validation process from [20]. This last one consists of a discriminator and a divergence estimator. The number of samples used from each distribution for each step is also highlighted: $N$ to generate samples, $M$ to train the discriminator, and $L$ to estimate the divergences.
@@ 3.3 Experimental design
In terms of the experimental design to evaluate the proposed method, and as described in the methodology section, we use the state-of-the-art VAE architecture [15] to train 10 different seeds. Subsequently, we apply methodologies based on transfer learning and meta-learning. For pre-training, we also include results using another state-of-the-art model, CTGAN. It is important to note that while we maintain the default parameters for CTGAN, we adjust the dimensionality of the latent space in the VAE depending on the dataset. This allows the VAE to capture the specific characteristics of each dataset more effectively. Particularly, we use a latent space dimension of $10$ for the Adult and Intrusion datasets, $20$ for the News dataset, and $15$ for the King dataset. We maintain a consistent hidden size of $256$ neurons for all VAE models. Regarding the configuration of parameters $M$ and $L$, we define two different validation configurations for each methodology: a reliable case and a more realistic case. The parameter $N$ remains unchanged, as it reflects the actual number of samples available to train the DGM, which is beyond our control. However, we can vary the number of generated samples for validation, i.e., $M$ and $L$. Specifically, the results presented for each dataset are as follows.
@@
  • “Big data’: First, we present the optimistic scenario with a sufficient $N$ of $10,000$ samples, where no methodology is needed to calculate the inductive bias. This provides results of the divergences that serve as an “upper bound" or reference for the best possible outcome. For $M$ and $L$, we maintain high values of the validation samples, $7500$ and $1000$, respectively.
@@
  • “Low data’: Next, we show results for a realistic scenario with few samples ($N=300$) without applying our methodology. This allows us to quantify the gain using the methodology and determine its benefits. For this case, we use two configurations for parameters $M$ and $L$: [$M=7500$, $L=1000$] and [$M=100$, $L=100$]. The second configuration is more realistic for few-data scenarios. When limited training data are available, there is also a limitation on the amount of data that can be effectively used for validation. The first configuration, with much larger values for $M$ and $L$, serves as a rigorous evaluation of the impact of our methodology. However, it is acknowledged that the use of small values for $M$ and $L$ can lead to unreliable metric estimations [20].
@@
  • “Pre-train’: In this case, we begin to apply the proposed methodology using the pre-training technique. Results are presented for both CTGAN and VAE. The parameter configurations chosen are: [$N=300$, $M=7500$, $L=1000$] and [$N=300$, $M=100$, $L=100$].
@@
  • “AVG’, “MAML’, “DRS’: These scenarios apply the model average (“AVG’) and the meta-learning techniques (“MAML’, “DRS’). We solely utilize the VAE architecture for multiple training runs. The following parameter configurations will be presented: [$N=300$, $M=7500$, $L=1000$] and [$N=300$, $M=100$, $L=100$].
@@
This setup aims to thoroughly evaluate the performance and robustness of the proposed methodologies in various data availability scenarios and parameter configurations.
@@ 3.4 Results
In this section, we present the results obtained from the experiments, focusing on scenarios that promote the validation of reliable synthetic data, characterized by higher values of $M$ and $L$. The results for scenarios with $M=100$ and $L=100$, considered unreliable due to their low information content, are presented in the Appendix for comparison purposes. For each database, we present a table summarizing the scenarios defined previously and their respective KL and JS divergence values. The results for each metric are displayed in the following format: mean (std) (lower is better). The code to replicate our results, along with the data used, can be found in our repository.
Scenario N VAE JS CTGAN JS VAE KL CTGAN KL
Big data 10000 0.079 (0.001) 0.150 (0.002) 0.153 (0.019) 0.420 (0.025)
Low data 300 0.331 (0.004) 0.563 (0.002) 0.697 (0.018) 1.653 (0.015)
[…]
AVG 300 0.157 (0.004) N/A 0.380 (0.043) N/A
MAML 300 0.300 (0.002) N/A 0.686 (0.037) N/A
DRS 300 0.189 (0.006) N/A 0.427 (0.043) N/A
Table 1: Adult dataset JS and KL results for each Scenario. Big data represents the ideal case where many samples ($N=10,000$) are available to generate reliable synthetic data. Low data represents a more realistic scenario in which a limited number of samples ($N=300$) are available, posing a challenge for synthetic data generation. The next rows compare the divergences obtained by each methodology (pre-training, model averaging, MAML, and DRS) applied to the low data scenario. Values in bold indicate an improvement due to the technique used. Results are represented as mean (std). Lower is better.
@@
Table 1 shows the validation results in terms of divergence obtained for the Adult dataset. We focus primarily on the JS divergence due to its interpretability as a bounded metric (ranging from 0 to 1). The table shows the upper and lower bounds used to assess the efficacy of the proposed methodology in the reliable case of higher validation samples ($M=7500$ and $L=1000$). These bounds are $0.079$ (upper) and $0.331$ (lower), highlighting a significant gap and room for improvement in the base VAE model (without any techniques applied). Examining the JS divergence results for the application of different proposed techniques, a consistent decrease in divergence is observed. The worst improvement is obtained for MAML ($0.300$) and the best for AVG ($0.157$). This implies that improvement is always present and, in the best cases, significantly high in terms of JS for VAE. A similar pattern is observed for KL divergence in VAE: better divergence results are obtained for transfer learning cases, but improvements are always achieved. However, for the CTGAN model (where only pre-training results are available), we see no significant improvement in either JS or KL divergence.
Table 2 summarizes the results obtained for the News dataset. The values demonstrate the effectiveness of the proposed techniques in improving divergence metrics compared to the established lower bounds. Notably, the VAE model shows improvement in divergence metrics in most methodologies, with the exception of the MAML technique. We hypothesize that the MAML technique might require a larger number of tasks to achieve comparable performance. Consistent with the findings for the previous dataset, Model Averaging emerges as the methodology that generally achieves the best results. When considering CTGAN, pre-training demonstrates improvement in the JS divergence metric. While the KL divergence results for pre-trained CTGAN do not show significant worsening compared to the lower bound, they remain within the established confidence intervals.
Scenario N VAE JS CTGAN JS VAE KL CTGAN KL
Big data 10000 0.253 (0.009) 0.463 (0.003) 0.647 (0.045) 1.506 (0.031)
Low data 300 0.840 (0.003) 0.962 (0.002) 4.582 (0.136) 8.994 (0.909)
[…]
AVG 300 0.609 (0.003) N/A 2.596 (0.060) N/A
MAML 300 0.851 (0.001) N/A 5.176 (0.242) N/A
DRS 300 0.645 (0.006) N/A 2.449 (0.057) N/A

### v2

@@ 2.2.2 DRS
Although MAML offers the potential to leverage the underlying structure of learning problems through a powerful optimization framework, it introduces a significant computational cost. Therefore, while we should seek a trade-off between accuracy and computational efficiency, there is no approach to managing this trade-off. An understanding of the domain-specific characteristics inherent to the meta-problem itself is needed.
DRS presents an alternative meta-learning approach that circumvents the computational burden of bilevel optimization problems. Unlike MAML, DRS trains a model on the combined data from all tasks. This eliminates the need for the complex optimization procedures present in MAML, leading to a more computationally efficient solution. However, it is important to acknowledge that DRS approximates the ideal solution (gao2020modeling, ).
@@
In this context, $\mathcal{D^{*}}$ represents the aggregated synthetic data collection, $x_{g}$, obtained across all training seeds $S$. We refer to this approach as Domain-Randomized Search due to its alignment with the domain randomization method presented in tobin2017domain and its core principle of directly searching over a distribution of domains (tasks).
Fig. 5 shows the application of the DRS procedure within the framework of our proposed methodology. Like the MAML case, $\theta_{DRS}$ is the initialization weights we aim to identify $\theta_{0}$.
Figure 5: Block diagram for the DRS case. The inductive bias is introduced by training a DGM $p_{\theta}$ on $x_{r}$ using $S$ different seeds. The synthetic dataset $x_{g}$ contains data generated by each seed and serves as input to DRS. The starting point for fine-tuning using the real data $x_{r}$ and obtaining $p_{\hat{\theta}}$ is the DRS solution, $\theta_{DRS}$.
Both MAML and DRS offer complementary approaches with a trade-off between modeling complexity and optimization cost (gao2020modeling, ). DRS delivers an approximate solution with lower computational demands, while MAML offers higher precision at the expense of greater computational resources. ‘DRS’ is also advantageous when dealing with a limited number of learning tasks. In our case, where data generated by each seed ($s=1,2,...,S$) is considered a task, and $S$ typically takes values around 101010, DRS is expected to provide better solutions than MAML, aligning with the findings of gao2020modeling . Finally, note that DRS is similar to the pre-training approach. While both techniques aim to improve model performance, they utilize data differently. Pre-training leverages data from the best VAE seed, whereas DRS capitalizes on data from all VAE seeds. This distinction reflects the core principle of DRS: exploring a wider range of possibilities by searching across a distribution of domains (tasks) represented by the various seed variations.
2.3 Summary and Comparison of Techniques
To clarify the overall methodology, we provide a high-level step-by-step summary of the artificial inductive bias generation process. (1) First, a DGM is trained using the limited real data $x_{r}$ or large synthetic data $x_{g}$ (sampled from a previously trained model), depending on the technique. (2) Then, an artificial inductive bias is constructed by either extracting weights (pre-training), averaging multiple trained models (model averaging), or applying a meta-learning algorithm across multiple training seeds (MAML or DRS). (3) These bias-derived weights $\theta_{0}$ are then used to initialize a second DGM, which is finally fine-tuned on the real data $x_{r}$ to obtain improved synthetic samples $\hat{x}_{g}$. This two-phase process enables the model to start learning from a more informative point in the parameter space, helping to mitigate overfitting or collapse in low-data regimes.
Each technique integrates this framework differently. Pre-training mitigates data scarcity by simulating a high-data regime with synthetically generated samples, while model averaging consolidates information from multiple VAE training seeds to create a more robust initialization. MAML optimizes an initialization that adapts quickly to new tasks, using the diversity of seed-generated data as a proxy for task variability. DRS simplifies this by combining all seeds into a single learning task, offering a trade-off between accuracy and computational burden. While pre-training is flexible and compatible with any architecture, it may transfer suboptimal representations if synthetic data quality is poor. Model averaging is simple and effective for VAEs, but it assumes that the mean of the parameter space corresponds to a good initialization. MAML can yield precise adaptations but may require more task diversity than available in low-seed settings. DRS is more robust under these conditions but offers a less principled optimization. Our approach addresses these limitations by leveraging the strengths of each technique within the VAE framework and evaluating them under consistent conditions. Limitations such as the architecture dependence of certain techniques and the impact of seed variability are discussed further in the results and conclusions.
@@ 3.1 Data
The experiments were conducted on three public datasets, one from the SDV environment (sdv, )–which also provides the CTGAN implementation used in this study–and the other two additional datasets from external repositories. We selected datasets with sufficient samples to allow multiple training and validation splits, ensuring a comprehensive evaluation of our method under different parameter settings.
  • Adult: The Adult Census Income dataset (adult, ) contains 32,561 samples with 14 mixed-type features (integer, categorical, binary). It is used to predict whether an individual’s annual income exceeds \$50,000\\$50,000\$ 50 , 000. The dataset has 13% missing values, concentrated in “work class" and “occupation."
  • King [^1] Source: https://www.kaggle.com/datasets/harlfoxem/housesalesprediction Accessed January $30^{\text{th}}$, 2025: The King County House Sales dataset contains 21,613 house sale records from King County, Washington, including Seattle, between May 2014 and May 2015. It includes 20 numerical and categorical features relevant to house pricing.
  • Letter: The Letter Recognition dataset (letter_recognition_59, ) includes 20,000 samples and 16 numerical features extracted from raster images of capital letters. The task is to classify each sample into one of 26 possible alphabet letters, making it a challenging multiclass classification problem. This dataset was obtained from the UCI Machine Learning Repository [^2] Source: https://archive.ics.uci.edu/dataset/59/letter+recognition Accessed June $17^{\text{th}}$, 2025.
This selection of datasets with varying sample sizes, feature dimensions, and task types (regression, binary, and multiclass classification) allows us to assess the performance and generalizability of the proposed method in diverse data scenarios. As will be demonstrated in the following sections, the proposed method is effective in generating synthetic data that retain the key characteristics of the original datasets across a wide range of applications. Additional datasets used for extended validation are described in B.
@@ 3.2 Validation metrics
To evaluate the effectiveness of our proposed method in capturing the real data distribution, we follow the validation approach from apellaniz2024synthetic . This method uses a probabilistic classifier (discriminator) to estimate the ratio of probability densities between real and synthetic data, enabling the calculation of KL and JS divergences. Unlike traditional validation methods that focus on individual features, this approach considers the entire data distribution, including complex relationships between features. Divergences provide robustness to noise and clear interpretability, making them ideal for assessing the similarity between real data $p(x_{r})$ and synthetic data $p_{\theta}$.
@@
The discriminator network is key to the validation process, trained to distinguish between real and synthetic data. It receives two sets of $M$ samples: one from the real data distribution $p(x_{r})$ (labeled as class 1) and another from the synthetic data distribution $p_{\theta}$ or $p_{\hat{\theta}}$ (labeled as class 0), depending on $N$, the dataset size. The discriminator learns a decision boundary during training, capturing differences between real and synthetic distributions. Once trained, it estimates KL and JS divergences by processing $L$ new samples from each distribution. The output probabilities of these samples are then used to compute the divergence metrics, providing a quantitative measure of distributional similarity.
Fig. 6 illustrates this process, highlighting the inductive bias generator, discriminator-based validation, and the different sample sizes used ($N$ for generation, $M$ for training, and $L$ for divergence estimation). As stated in apellaniz2024synthetic , $M$ and $L$ must be large enough to ensure reliable divergence estimates, which we consider in our experiments.
@@
Figure 6: General Scheme of the proposed approach. Overall scheme of the approach and its validation process from apellaniz2024synthetic . This last one consists of a discriminator and a divergence estimator. The number of samples used from each distribution for each step is also highlighted: $N$ to generate samples, $M$ to train the discriminator, and $L$ to estimate the divergences.
In addition to this divergence-based evaluation, we incorporate a utility-based validation protocol (also referred to as machine learning efficacy in the literature (e.g., xu2019modeling ). This approach assesses the utility of synthetic data in downstream predictive tasks by training machine learning models with synthetic samples and evaluating their performance on real data. Specifically, we define two configurations:
  • Real metric: The model is trained and validated using real data. This provides an upper-bound reference for predictive performance.
  • Synthetic metric: The model is trained on synthetic data and evaluated on real data. This reflects how well the synthetic data preserve task-relevant structure.
Higher classification metrics or lower regression errors in the Synthetic configuration indicate that the generated data maintain high utility for downstream tasks. In classification tasks, the utility metric corresponds to accuracy, while in regression tasks it corresponds to the MSE. Together, both divergence-based and utility-based evaluations provide a more comprehensive view of the quality and applicability of the synthetic data.
@@ 3.3 Experimental design
To evaluate the proposed methodology, we designed a comprehensive experimental pipeline including both divergence-based and utility-based validation. As described in the methodology section, we use the state-of-the-art VAE architecture (apellaniz2024improved, ) and train 10 different seeds per experiment to account for stochastic variability and to enable the application of meta-learning and averaging techniques. We also evaluate the pre-training strategy using CTGAN, a well-established conditional GAN model for generating tabular data.
The VAE used in our experiments is based on a VAE with a BGM (apellaniz2024improved, ), which replaces the standard Gaussian latent assumption with a learned mixture model, enabling more flexible and expressive latent representations. Both encoder and decoder are fully connected neural networks with ReLU activations, a hidden layer of 256256256 neurons, and dataset-specific latent dimensions: 101010 for Adult, Intrusion, Letter, and PenDigits; 151515 for King; and 202020 for News. The model is trained using the standard ELBO loss, which includes a KL divergence regularizer and employs early stopping to prevent overfitting. All training is conducted on the CPU with a batch size of 256256256 and a default learning rate of $1\times 10^{-3}$.
The CTGAN model (xu2019modeling, ) is implemented using the SDV library’s default configuration (sdv, ), with 1,0001,0001 , 000 training epochs and GPU acceleration enabled. It employs a conditional generator and a discriminator trained adversarially and includes a conditional vector strategy to handle categorical imbalance and mode-specific normalization for numerical features.
We use 5-fold cross-validation throughout all experiments, consistently splitting the real data into training and validation folds. For both divergence and utility validations, we define two configurations: a reliable setting with $M=7500$ and $L=1000$ samples and a more realistic low-resource setting with $M=L=100$. The sample size $N$ for training the generative model is fixed based on the scenario being simulated and is not manipulated.
Utility-based evaluation is conducted across classification and regression tasks, depending on the dataset. In each case:
  • Classification uses a custom neural network with three hidden layers (256–64–32), batch normalization, dropout, and LeakyReLU activations. Training runs for 10,00010,00010 , 000 epochs with early stopping.
  • Regression employs a single-layer linear model trained for 2,0002,0002 , 000 epochs using Mean Squared Error (MSE) loss and Adam optimizer.
The experimental results are grouped into the following scenarios:
  1. 1.
@@
    ‘Big data’: First, we present the optimistic scenario with a sufficient $N$ of 10,00010,00010 , 000 samples, where no methodology is needed to calculate the inductive bias. While we acknowledge that the term ‘big data’ is typically associated with datasets containing hundreds of thousands or even millions of records, in this experimental design, we use it in a relative sense to contrast this setting with the data-scarce scenarios that are the primary focus of our study. This configuration provides divergence results that serve as an “upper bound" or reference for the best possible outcome. For $M$ and $L$, we maintain high values of the validation samples, 750075007500 and 100010001000, respectively.
  2. 2.
@@
    ‘Low data’: Next, we show results for a realistic scenario with few samples ($N=300$) without applying our methodology. This allows us to quantify the gain using the method and determine its benefits. For this case, we use two configurations for parameters $M$ and $L$: [$M=7500$, $L=1000$] and [$M=100$, $L=100$]. The second configuration is more realistic for few-data scenarios. When limited training data are available, there is also a limitation on the amount of data that can be effectively used for validation. The first configuration, with much larger values for $M$ and $L$, serves as a rigorous evaluation of the impact of our methodology. However, it is acknowledged that using small values for $M$ and $L$ can lead to unreliable metric estimations (apellaniz2024synthetic, ).
  3. 3.
@@
    ‘Pre-train’: In this case, we apply the proposed methodology using the pre-training technique. Results are presented for both CTGAN and VAE. The parameter configurations chosen are: [$N=300$, $M=7500$, $L=1000$] and [$N=300$, $M=100$, $L=100$].
  4. 4.
@@
    ‘AVG,’ ‘MAML,’ ‘DRS’: These scenarios apply the model average (‘AVG’) and the meta-learning techniques (‘MAML,’ ‘DRS’). We solely utilize the VAE architecture for multiple training runs. The following parameter configurations will be presented: [$N=300$, $M=7500$, $L=1000$] and [$N=300$, $M=100$, $L=100$].
@@
This setup aims to thoroughly evaluate the performance and robustness of the proposed methodologies in various data availability scenarios and parameter configurations. Hardware and software specifications are detailed in C.
@@ 3.4 Results
In this section, we present the experimental results, focusing on scenarios with higher values of $M$ and $L$, which promote more reliable divergence estimation. Additional results for other datasets and the low-resource configuration ($M=100$, $L=100$) are provided in A for completeness. For each database, we summarize the outcomes across all scenarios using both divergence-based and utility-based validation protocols. Divergence metrics include JS and KL divergences. Utility metrics reflect the performance of a downstream model trained with synthetic data. Specifically, we consider two setups: training and validating with real data (Real metric) and training with synthetic data and validating on real data (Synth metric). For classification tasks, the metric used is accuracy; for regression tasks, the metric used is the MSE. All results are reported as mean (standard deviation) across 10 independent runs with different random seeds.
The code to replicate our results, along with the data used, can be found in our repository. Additionally, C presents a comparative summary of the computational costs associated with each method, including training times and hardware and software configurations. This analysis supports practitioners in evaluating trade-offs between performance and resource requirements when selecting a technique.
@@
Subtable (a) in Table˜1 shows the validation results obtained for the Adult dataset. Regarding the divergence estimations, we focus primarily on the JS divergence due to its interpretability as a bounded metric (ranging from 0 to 1). The table shows the upper and lower bounds used to assess the efficacy of the proposed methodology in the reliable case of higher validation samples ($M=7500$ and $L=1000$). These bounds are 0.1150.1150.115 (upper) and 0.4200.4200.420 (lower), highlighting a significant gap and room for improvement in the base VAE model (without any techniques applied). A consistent decrease in divergence is observed when examining the JS divergence results for the application of different proposed techniques. The worst improvement is obtained for ‘MAML’ (0.3580.3580.358) and the best for ‘AVG’ (0.1630.1630.163). This implies that improvement is always present and, in the best cases, significantly high in JS for VAE. A similar pattern is observed for KL divergence in VAE: better divergence results are obtained for transfer learning cases, but improvements are always achieved. For the CTGAN model (where only pre-training results are available), we also observe a significant improvement in both JS and KL divergence despite the generation process appearing to be less accurate in terms of divergence estimation compared to the VAE. Regarding the utility validation, training with synthetic data generated by the VAE in the ‘Big data’ scenario yields accuracy levels close to those obtained with real data. In the ‘Low data’ case, the accuracy drops when using synthetic data but increases again when any of the proposed techniques are applied, validating the utility of these strategies. However, no improvement is observed in utility metrics when applying the CTGAN-based pre-training, suggesting that while divergence metrics may improve, they do not necessarily translate into enhanced downstream performance for this model.
Subtable (b) in Table˜1 further reinforces the efficacy of the proposed methodologies on the King dataset. The VAE model consistently improves on the lower bounds established for the JS divergence across all techniques. In terms of KL divergence, the most notable improvement is observed with the ‘AVG’ technique, which significantly reduces the KL value from 5.9635.9635.963 to 3.3223.3223.322. In contrast, CTGAN results are less favorable, with divergences that remain close to or above the baseline (‘Low data’) and offer little gain in the presence of pre-training. These results suggest that CTGAN is less robust in modeling the complex structure of regression data in small-sample regimes. The utility validation, as measured by MSE, mirrors the trends observed in the divergence results for both generative models. The synthetic data generated by the VAE under ‘Low data’ conditions shows a substantial increase in MSE compared to the real-data baseline, indicating reduced predictive utility. However, applying techniques such as ‘AVG,’ ‘Pre-train,’ or ‘DRS’ significantly reduces the MSE, bringing it closer to the values obtained in the ‘Big data’ scenario. This validates the practical benefits of introducing inductive biases even in regression contexts. The technique ‘MAML,’ in this case, appears less effective in terms of utility, likely due to the higher sensitivity of regression tasks to noise and instability during meta-learning optimization.
Ultimately, subtable (c) in Table˜1 presents the results obtained for the Letter dataset, which involves a challenging multiclass classification task with 26 output classes. Regarding the divergence metrics, the application of the proposed techniques consistently improves JS divergence for the VAE model, except for ‘MAML,’ which yields no improvement over the baseline scenario. A similar trend is observed in KL divergence: while most techniques achieve a notable reduction, ‘MAML’ again fails to improve over the baseline, possibly due to optimization instability or suboptimal adaptation in this high-class-cardinality setting. For the CTGAN model, we observe no improvement in JS divergence but a significant reduction in KL divergence. In terms of utility validation, measured via accuracy, the only meaningful improvement over the ‘Low data’ baseline is observed with the ‘DRS’ technique. All other methods yield marginal gains or similar performance, indicating that despite improvements in distributional similarity, they do not translate into enhanced classification performance for this dataset. This may be due to the complexity of the classification task or the need for more targeted biasing strategies to capture high-granularity class structures better.
Scenario N DIVERGENCE VALIDATION UTILITY VALIDATION
VAE JS CTGAN JS VAE KL CTGAN KL Real Acc VAE Synth Acc CTGAN Synth Acc
Big data 10000 0.115 (0.002) 0.136 (0.002) 0.272 (0.014) 0.385 (0.025) 0.777 (0.001) 0.748 (0.002) 0.739 (0.006)
Low data 300 0.420 (0.003) 0.742 (0.003) 1.014 (0.036) 2.454 (0.027) 0.775 (0.005) 0.673 (0.006) 0.682 (0.014)
Pre-train 300 0.182 (0.002) 0.565 (0.002) 0.413 (0.008) 1.665 (0.025) N/A 0.731 (0.005) 0.689 (0.003)
AVG 300 0.163 (0.003) N/A 0.391 (0.030) N/A N/A 0.728 (0.004) N/A
MAML 300 0.358 (0.004) N/A 0.876 (0.041) N/A N/A 0.702 (0.006) N/A
DRS 300 0.180 (0.002) N/A 0.463 (0.037) N/A N/A 0.721 (0.007) N/A
(a) Adult dataset
Scenario N DIVERGENCE VALIDATION UTILITY VALIDATION
VAE JS CTGAN JS VAE KL CTGAN KL Real Acc VAE Synth MSE CTGAN Synth MSE
Big data 10000 0.742 (0.005) 0.746 (0.004) 3.243 (0.152) 2.920 (0.099) 0.325 (0.005) 0.383 (0.020) 0.369 (0.001)
Low data 300 0.919 (0.002) 0.916 (0.003) 5.963 (0.135) 6.751 (0.394) 0.329 (0.006) 0.516 (0.011) 0.756 (0.029)
Pre-train 300 0.877 (0.002) 0.948 (0.003) 6.215 (0.350) 8.394 (0.328) N/A 0.451 (0.012) 0.767 (0.015)
AVG 300 0.724 (0.002) N/A 3.322 (0.235) N/A N/A 0.366 (0.012) N/A
MAML 300 0.899 (0.002) N/A 10.912 (0.472) N/A N/A 0.686 (0.065) N/A
DRS 300 0.871 (0.007) N/A 6.118 (0.710) N/A N/A 0.457 (0.025) N/A
(b) King dataset
Scenario N DIVERGENCE VALIDATION UTILITY VALIDATION
VAE JS CTGAN JS VAE KL CTGAN KL Real Acc VAE Synth Acc CTGAN Synth Acc
Big data 10000 0.395 (0.004) 0.635 (0.005) 0.911 (0.027) 1.678 (0.059) 0.815 (0.007) 0.651 (0.005) 0.075 (0.006)
Low data 300 0.654 (0.007) 0.977 (0.003) 1.988 (0.029) 10.965 (0.705) 0.815 (0.005) 0.410 (0.008) 0.039 (0.010)
Pre-train 300 0.585 (0.005) 0.978 (0.001) 1.694 (0.035) 8.728 (0.650) N/A 0.394 (0.010) 0.032 (0.006)
AVG 300 0.581 (0.007) N/A 1.744 (0.063) N/A N/A 0.385 (0.005) N/A
MAML 300 0.669 (0.005) N/A 2.001 (0.056) N/A N/A 0.318 (0.010) N/A
DRS 300 0.608 (0.005) N/A 1.635 (0.061) N/A N/A 0.437 (0.012) N/A
(c) Letter dataset